Paula is 12 years older than brother Bob. In 4 years, Bob will be 2/3 as old as Paula. How old are they now?
Answer by jermz69(5)

Paula is 12 years older than brother Bob. In 4 years, Bob will be 2/3 as old as Paula. How old are they now?

Let x: Bob's age
x + 12: Paula's age

In four years:
Bob's age: x + 4
Paula's age: x + 16

Equation: x+4=(2/3)(x + 16)
3(x+4)=(3)(2/3)(x+16)
3x+12=2(x+16)
3x+12=2x+32
3x-2x=32-12
x=20

20+12=32

Bob's age now: 20
Paula's age: 32

Check: x+4=(2/3)(x + 16)
20+4=(2/3)(20+16)
24=(2/3)(36)
24=24

By the way, I'm jermz (my nickname). Hope I helped!